OpenStudy (anonymous):

Simplify the expression completely:(3 y^8)^3/y^(- 7)

jigglypuff314 (jigglypuff314):

first, you can bring the y^(-7) up so \[\frac{ (3y ^{8})^{3} }{ y ^{-7} } \rightarrow (3y ^{8})^{3}\times y ^{7}\] then you can distribute the exponent into the parenthesis so (3y^8)^3 = 27y^24 so (27y^24)(y^7)=27y^31

OpenStudy (anonymous):

Thanks! It make complete sense now :)